OpenStudy (reverdeparadis):

Solve using elimination. 7x - 3y = 20 5x + 3y = 16 Is it (3, 1/3)?

OpenStudy (jdoe0001):

is that what you got ?

OpenStudy (jdoe0001):

what did you eliminate anyway?

OpenStudy (ybarrap):

Let's plug in, $$ 7\times\color{red}{ 3}-\cfrac{3}{\color{red}{3}}=21-1=20~\checkmark\\ 5\times \color{red}{3} +\cfrac{3}{\color{red}{3}}=15+1=16~\checkmark\\ $$ Looks good!